# STHN: Deep Homography Estimation for UAV Thermal Geo-localization with Satellite Imagery

### v2

@@ Abstract
Index Terms:
Deep Learning for Visual Perception, Aerial Systems: Applications, Localization
@@ I Introduction
Addressing these limitations, this study introduces Satellite-Thermal Homography Network (STHN) framework (see Fig. 1) that leverages deep homography estimation techniques [12, 13, 14, 15, 16, 17, 18] to directly align thermal images with satellite maps of the local region, optimizing localization in GNSS-denied scenarios. This approach adopts a two-stage coarse-to-fine strategy: 1) Coarse alignment, which matches small thermal images to large satellite maps within a search radius of $512~{}$\mathrm{m}$$ with a challenging constant size ratio of $11\%$; and 2) Refinement, which crops and resizes the selected region and applies a second-stage estimation for enhanced accuracy.
@@ IV-A Dataset
For training and evaluation, our study utilizes the Boson-nighttime [11] real-world dataset which contains $10,256$ train pairs, $13,011$ validation pairs, and $26,568$ test pairs of coupled satellite RGB and nadir-view 8-bit thermal imagery. We have expanded the dataset by augmenting the collection of satellite images without corresponding thermal images from $79,950$ to $163,344$ images, covering an area of $215.78~{}$\mathrm{k}\mathrm{m}$^{2}$. This enhancement focuses on the desert and farm areas near the original dataset’s sampling region, thereby incorporating a broader spectrum of geographical patterns. Additionally, the test region is then excluded from the generated data to ensure a robust evaluation of generalization performance. The thermal images in the dataset are captured between 9:00 PM and 4:00 AM, and they are aligned with an approx. spatial resolution of $1~{}$\mathrm{m}\mathrm{/}\mathrm{p}\mathrm{x}$$. The thermal images are cropped to $W_{T}\times W_{T}$ pixels ($\mathrm{p}\mathrm{x}$), where $W_{T}=512$. The satellite images  [^1] Bing RGB satellite imagery is sourced from Maxar: https://www.bing.com/maps/aerial are cropped to $W_{S}\times W_{S}$. Fig. 3 shows the ground truth overlap between thermal images and satellite images with different $W_{S}$. For $W_{S}=512/1024/1536$, the size ratios between thermal images and satellite images are $100\%$, $25\%$, and $11\%$.
@@ IV-C Implementation Details
For pre-processing, the resize side length $W_{R}$ is $256~{}$\mathrm{p}\mathrm{x}$$. The training iteration numbers of the coarse alignment and refinement modules are $200000$ with a batch size of $16$. The AdamW optimizer [37] is employed for model training, utilizing a linear learning rate decay scheduler with warmup with the peak learning rate at $1\mathrm{e}{-4}$. The numbers of iterative updates $K_{1}$ and $K_{2}$ are both set to $6$. Depending on the setting, the correlation module’s level is $2$ (for $W_{S}=512$) or $4$ (for $W_{S}=1024,1536$) with a search radius of $4$. For bounding box augmentation, $\Delta p_{1},\Delta p_{2}$ is set to vary between $(-64,64)$, $\Delta p_{3}$ is set within $[0,64)$, and $\Delta p_{4}$ is $64$ by parameter tuning. For geometric noises, we extend the coverage of thermal images, apply corresponding data augmentations, and center crop the thermal images to avoid black padding on their boundary. Our models are developed using PyTorch. The inference speed is measured with one NVIDIA RTX-2080-Ti GPU.
@@ V Results
Methods $D_{C}=50~{}$\mathrm{m}$$ $D_{C}=64~{}$\mathrm{m}$$ $D_{C}=128~{}$\mathrm{m}$$ $D_{C}=256~{}$\mathrm{m}$$ $D_{C}=512~{}$\mathrm{m}$$ Failure Rate
@@
TABLE II: Comparison between different image-based matching methods and our estimation method when $D_{C}=512~{}$\mathrm{m}$$.
@@ V-A Comparison with Baselines
Subsequently, our analysis compares our methods with various deep homography estimation frameworks, including DHN [13], LocalTrans [16], and IHN [12] (state-of-the-art method in real-time applications). These baselines with one-stage models are trained on the Boson-nighttime dataset. We report the baseline results considering $W_{S}=512$ as representative results since other $W_{S}$ show similar trends in our analysis. The results show the superior performance of our approach for satellite-thermal alignment and geo-localization. A notable observation from the data is the different performance preferences across varying $D_{C}$ distances: for $D_{C}=50~{}$\mathrm{m}$$ and $D_{C}=64~{}$\mathrm{m}$$, the optimal $W_{S}$ is $512$, while for mid-range distances of $D_{C}=128~{}$\mathrm{m}$$ and $D_{C}=256~{}$\mathrm{m}$$, using $W_{S}=1024$ leads to the best results. Additionally, for the longest distance of $D_{C}=512~{}$\mathrm{m}$$, our novel two-stage method with $W_{S}=1536$ emerges as the most effective strategy. The findings indicate that for cases where $D_{C}\leq 256~{}$\mathrm{m}$$, employing our one-stage method combined with a carefully chosen $W_{S}$ emerges as the most effective strategy. Further explanation of the correlation between $W_{S}$ and $D_{C}$ is in Section V-B.
We find that our two-stage method fails to enhance performance for distances $D_{C}=50~{}$\mathrm{m}$$, $64~{}$\mathrm{m}$$, and $128~{}$\mathrm{m}$$, instead leading to a decline in accuracy. Upon examining the visualized outcomes, we observe that for smaller distances ($D_{C}\leq 128~{}$\mathrm{m}$$), the initial coarse alignment is sufficiently accurate, making the refinement module’s excessive iterative updates introduce noise into the final predictions, thereby degrading performances. Nevertheless, our two-stage approach maintains an overall MACE of less than $15~{}$\mathrm{m}$$ across all considered $D_{C}$, establishing robust baselines for this task. Notably, for achieving precise geo-localization at $D_{C}=512~{}$\mathrm{m}$$, this two-stage strategy demonstrates the best performance, underscoring its effectiveness for large-scale search regions.
@@ V-B2 Coarse alignment
Fig. 5 illustrates the correlation between validation MACE and $W_{S}$ across various $D_{C}$. The figure shows that as $W_{S}$ increases, the validation MACE for smaller translation distances ($D_{C}=50~{}$\mathrm{m}$$ and $64~{}$\mathrm{m}$$) slightly increases, suggesting a deterioration in alignment accuracy. In contrast, for larger translation distances ($D_{C}=128~{}$\mathrm{m}$$, $256~{}$\mathrm{m}$$, $512~{}$\mathrm{m}$$), the validation MACE decreases, indicating improved alignment accuracy. The intuition is that an increase in $W_{S}$, without a corresponding adjustment in $W_{R}$, leads to a higher pixel-per-meter (ppm) ratio after image resizing. This increment in ppm ratio can negatively affect the alignment accuracy. Conversely, a larger $W_{S}$ enhances alignment accuracy for greater translation ($D_{C}$), especially for $W_{S}=1536$ and $D_{C}=512~{}$\mathrm{m}$$. In these cases, a larger $W_{S}$ ensures the full coverage of the thermal image, which is crucial for accurately calculating correlation volumes $\mathbf{C}$.
@@ V-C Robustness Evaluation and Visualization
Ideally, the UAV onboard compass and gimbal camera would supply precise data, enabling the accurate alignment of images to the north. However, it is crucial for our algorithm to demonstrate tolerance towards certain rotation and perspective transformation inaccuracies during active flights. Additionally, understanding how our algorithm performs when there is a change in flight altitude—which results in a change of the thermal image’s coverage area, denoted as resizing noise—is essential. To assess the algorithm’s robustness under these conditions, we perform experiments that introduce specific rotation, resizing, and perspective transformation noises. For rotation disturbances, the thermal images undergo random rotations up to $5^{\circ}$, $10^{\circ}$, or $30^{\circ}$. For resizing disturbances, the images are randomly scaled by a factor of $1+\Delta r$, with $\Delta r$ varying within either ±$\pm 0.1$, ±$\pm 0.2$, or ±$\pm 0.3$. For perspective transformation, we randomly adjust the four corners of $512\times 512$ thermal images up to $8~{}$\mathrm{p}\mathrm{x}$$, $16~{}$\mathrm{p}\mathrm{x}$$, or $32~{}$\mathrm{p}\mathrm{x}$$.
In Table III, we evaluate the robustness of our one-stage and two-stage strategies against a variety of geometric noise conditions with $D_{C}=512~{}$\mathrm{m}$$ and $W_{S}=1536$. The analysis indicates a significant decrease in performance for the one-stage method under these conditions, in contrast to the two-stage strategy, which demonstrates a notable robustness against geometric perturbations. Specifically, the two-stage strategy effectively maintains test MACE below $22~{}$\mathrm{m}$$ and test CE below $20~{}$\mathrm{m}$$ in most scenarios, with notable exceptions being in instances of $30^{\circ}$ rotation noise. While incremental perspective transformations and resizing have minimal impact on accuracy, large rotation noise can significantly degrade performance. This suggests the tolerance of our strategies to different types of geometric noise. Overall, the results validate our method’s robustness and its ability to estimate these disturbances, underscoring the two-stage strategy’s superior effectiveness and reliability in mitigating the negative effects of these disturbances. Fig. 7 further illustrates this point by showcasing visual comparisons between the failure instances of the one-stage method and the success cases of the two-stage method, demonstrating the latter’s improved robustness.
@@ VI Conclusions
$8~{}$\mathrm{p}\mathrm{x}$$ 30.53 26.41 13.70 10.94
$16~{}$\mathrm{p}\mathrm{x}$$ 34.07 27.22 16.24 11.73
$32~{}$\mathrm{p}\mathrm{x}$$ 44.03 30.35 20.90 14.11

### v3

@@ I Introduction
Addressing these limitations, this study introduces Satellite-Thermal Homography Network (STHN) framework (see Fig. 1) that leverages deep homography estimation techniques [12, 13, 14, 15, 16, 17, 18] to directly align thermal images with satellite maps of the local region, optimizing localization in GNSS-denied scenarios. This approach adopts a two-stage coarse-to-fine strategy: 1) Coarse alignment, which matches small thermal images to large satellite maps within a search radius of $512~$\mathrm{m}$\$ with a challenging constant size ratio of $11\%$; and 2) Refinement, which crops and resizes the selected region and applies a second-stage estimation for enhanced accuracy.
@@ IV-A Dataset
For training and evaluation, our study utilizes the Boson-nighttime [11] real-world dataset which contains $10,256$ train pairs, $13,011$ validation pairs, and $26,568$ test pairs of coupled satellite RGB and nadir-view 8-bit thermal imagery. We have expanded the dataset by augmenting the collection of satellite images without corresponding thermal images from $79,950$ to $163,344$ images, covering an area of $215.78~$\mathrm{k}\mathrm{m}$^{2}$. This enhancement focuses on the desert and farm areas near the original dataset’s sampling region, thereby incorporating a broader spectrum of geographical patterns. Additionally, the test region is then excluded from the generated data to ensure a robust evaluation of generalization performance. The thermal images in the dataset are captured between 9:00 PM and 4:00 AM, and they are aligned with an approx. spatial resolution of $1~$\mathrm{m}\mathrm{/}\mathrm{p}\mathrm{x}$\$. The thermal images are cropped to $W_{T}\times W_{T}$ pixels ($\mathrm{p}\mathrm{x}$), where $W_{T}=512$. The satellite images  [^1] Bing RGB satellite imagery is sourced from Maxar: https://www.bing.com/maps/aerial are cropped to $W_{S}\times W_{S}$. Fig. 3 shows the ground truth overlap between thermal images and satellite images with different $W_{S}$. For $W_{S}=512/1024/1536$, the size ratios between thermal images and satellite images are $100\%$, $25\%$, and $11\%$.
@@ IV-C Implementation Details
For pre-processing, the resize side length $W_{R}$ is $256~$\mathrm{p}\mathrm{x}$\$. The training iteration numbers of the coarse alignment and refinement modules are $200000$ with a batch size of $16$. The AdamW optimizer [37] is employed for model training, utilizing a linear learning rate decay scheduler with warmup with the peak learning rate at $1\mathrm{e}{-4}$. The numbers of iterative updates $K_{1}$ and $K_{2}$ are both set to $6$. Depending on the setting, the correlation module’s level is $2$ (for $W_{S}=512$) or $4$ (for $W_{S}=1024,1536$) with a search radius of $4$. For bounding box augmentation, $\Delta p_{1},\Delta p_{2}$ is set to vary between (64,64)(-64,64), $\Delta p_{3}$ is set within $[0,64)$, and $\Delta p_{4}$ is $64$ by parameter tuning. For geometric noises, we extend the coverage of thermal images, apply corresponding data augmentations, and center crop the thermal images to avoid black padding on their boundary. Our models are developed using PyTorch. The inference speed is measured with one NVIDIA RTX-2080-Ti GPU.
@@ V Results
Methods $D_{C}=50~$\mathrm{m}$\$ $D_{C}=64~$\mathrm{m}$\$ $D_{C}=128~$\mathrm{m}$\$ $D_{C}=256~$\mathrm{m}$\$ $D_{C}=512~$\mathrm{m}$\$ Failure Rate
@@
TABLE II: Comparison between different image-based matching methods and our estimation method when $D_{C}=512~$\mathrm{m}$\$.
@@ V-A Comparison with Baselines
Subsequently, our analysis compares our methods with various deep homography estimation frameworks, including DHN [13], LocalTrans [16], and IHN [12] (state-of-the-art method in real-time applications). These baselines with one-stage models are trained on the Boson-nighttime dataset. We report the baseline results considering $W_{S}=512$ as representative results since other $W_{S}$ show similar trends in our analysis. The results show the superior performance of our approach for satellite-thermal alignment and geo-localization. A notable observation from the data is the different performance preferences across varying $D_{C}$ distances: for $D_{C}=50~$\mathrm{m}$\$ and $D_{C}=64~$\mathrm{m}$\$, the optimal $W_{S}$ is $512$, while for mid-range distances of $D_{C}=128~$\mathrm{m}$\$ and $D_{C}=256~$\mathrm{m}$\$, using $W_{S}=1024$ leads to the best results. Additionally, for the longest distance of $D_{C}=512~$\mathrm{m}$\$, our novel two-stage method with $W_{S}=1536$ emerges as the most effective strategy. The findings indicate that for cases where $D_{C}\leq 256~$\mathrm{m}$\$, employing our one-stage method combined with a carefully chosen $W_{S}$ emerges as the most effective strategy. Further explanation of the correlation between $W_{S}$ and $D_{C}$ is in Section V-B.
We find that our two-stage method fails to enhance performance for distances $D_{C}=50~$\mathrm{m}$\$, $64~$\mathrm{m}$\$, and $128~$\mathrm{m}$\$, instead leading to a decline in accuracy. Upon examining the visualized outcomes, we observe that for smaller distances ($D_{C}\leq 128~$\mathrm{m}$\$), the initial coarse alignment is sufficiently accurate, making the refinement module’s excessive iterative updates introduce noise into the final predictions, thereby degrading performances. Nevertheless, our two-stage approach maintains an overall MACE of less than $15~$\mathrm{m}$\$ across all considered $D_{C}$, establishing robust baselines for this task. Notably, for achieving precise geo-localization at $D_{C}=512~$\mathrm{m}$\$, this two-stage strategy demonstrates the best performance, underscoring its effectiveness for large-scale search regions.
@@ V-B2 Coarse alignment
Fig. 5 illustrates the correlation between validation MACE and $W_{S}$ across various $D_{C}$. The figure shows that as $W_{S}$ increases, the validation MACE for smaller translation distances ($D_{C}=50~$\mathrm{m}$\$ and $64~$\mathrm{m}$\$) slightly increases, suggesting a deterioration in alignment accuracy. In contrast, for larger translation distances ($D_{C}=128~$\mathrm{m}$\$, $256~$\mathrm{m}$\$, $512~$\mathrm{m}$\$), the validation MACE decreases, indicating improved alignment accuracy. The intuition is that an increase in $W_{S}$, without a corresponding adjustment in $W_{R}$, leads to a higher pixel-per-meter (ppm) ratio after image resizing. This increment in ppm ratio can negatively affect the alignment accuracy. Conversely, a larger $W_{S}$ enhances alignment accuracy for greater translation ($D_{C}$), especially for $W_{S}=1536$ and $D_{C}=512~$\mathrm{m}$\$. In these cases, a larger $W_{S}$ ensures the full coverage of the thermal image, which is crucial for accurately calculating correlation volumes $\mathbf{C}$.
@@ V-C Robustness Evaluation and Visualization
Ideally, the UAV onboard compass and gimbal camera would supply precise data, enabling the accurate alignment of images to the north. However, it is crucial for our algorithm to demonstrate tolerance towards certain rotation and perspective transformation inaccuracies during active flights. Additionally, understanding how our algorithm performs when there is a change in flight altitude—which results in a change of the thermal image’s coverage area, denoted as resizing noise—is essential. To assess the algorithm’s robustness under these conditions, we perform experiments that introduce specific rotation, resizing, and perspective transformation noises. For rotation disturbances, the thermal images undergo random rotations up to $5^{\circ}$, $10^{\circ}$, or $30^{\circ}$. For resizing disturbances, the images are randomly scaled by a factor of $1+\Delta r$, with $\Delta r$ varying within either $\pm 0.1$, $\pm 0.2$, or $\pm 0.3$. For perspective transformation, we randomly adjust the four corners of $512\times 512$ thermal images up to $8~$\mathrm{p}\mathrm{x}$\$, $16~$\mathrm{p}\mathrm{x}$\$, or $32~$\mathrm{p}\mathrm{x}$\$.
In Table III, we evaluate the robustness of our one-stage and two-stage strategies against a variety of geometric noise conditions with $D_{C}=512~$\mathrm{m}$\$ and $W_{S}=1536$. The analysis indicates a significant decrease in performance for the one-stage method under these conditions, in contrast to the two-stage strategy, which demonstrates a notable robustness against geometric perturbations. Specifically, the two-stage strategy effectively maintains test MACE below $22~$\mathrm{m}$\$ and test CE below $20~$\mathrm{m}$\$ in most scenarios, with notable exceptions being in instances of $30^{\circ}$ rotation noise. While incremental perspective transformations and resizing have minimal impact on accuracy, large rotation noise can significantly degrade performance. This suggests the tolerance of our strategies to different types of geometric noise. Overall, the results validate our method’s robustness and its ability to estimate these disturbances, underscoring the two-stage strategy’s superior effectiveness and reliability in mitigating the negative effects of these disturbances. Fig. 7 further illustrates this point by showcasing visual comparisons between the failure instances of the one-stage method and the success cases of the two-stage method, demonstrating the latter’s improved robustness.
@@ VI Conclusions
$8~$\mathrm{p}\mathrm{x}$\$ 30.53 26.41 13.70 10.94
$16~$\mathrm{p}\mathrm{x}$\$ 34.07 27.22 16.24 11.73
$32~$\mathrm{p}\mathrm{x}$\$ 44.03 30.35 20.90 14.11